OpenStudy (anonymous):

The measurement of the height of 600 students of a college is normally distributed with a mean of 175 centimeters and a standard deviation of 5 centimeters. What percent of students are less than 170 cm in height?

OpenStudy (amistre64):

youll need a modified zscore ... scale it by sqrt(n)

OpenStudy (amistre64):

\[z=\frac{x-\bar x}{\sigma}\] \[z\sqrt n=\frac{x-\bar x}{\sigma}\sqrt n\] \[z'=\frac{x-\bar x}{\sigma/\sqrt n}\]